hi im 40 and in algebra at my college and im stuck please help!!!
(2/5)^2*3/8+1/5 divided by 3/4

Answer by Maths68(1474)

(2/5)^2*3/8+1/5 divided by 3/4
Use BODMAS rule
First solve Brackets
then Order (powers etc.)
then Divide
then Multiply
then Add
then Subtract
here we go
=[(2/5)^2*3/8+1/5]/3/4
=[(4/25)*3/8 + 1/5]/3/4
=[(4*3/25*8)+1/5]/3/4
=[12/200+1/5]/3/4
=[3/50+1/5]/3/4
Taking LCM
=[(3+10)/50]/3/4
=(13/50)/3/4 (inverse 3/4 and change the divide sign to multiply)
=(13/50) * 4/3
=(13*4) / (50 * 3)
=(13*2) / (25 * 3)
=26/75

=